# Xing Huang(a), Eva Kopfer(b), Panpan Ren(c) a)

### v3

@@ 2.2 Log-Sobolev inequality for the invariant probability measure of time-homogeneous decoupled SDEs
Before we prove our next main result Theorem 2.6, we make a few comments on the proof. First, by Wang’s Harnack inequality (2.2) for $\tilde{P}_{t}^{\mu}$, the hyperboundedness (2.26) for $\tilde{P}_{t}^{\mu}$ holds. From this we can employ Theorem 2.1(1) in [28], which provides the defective log-Sobolev inequality for the invariant probability measure (IPM) of (2.13). Let us remark, that although [28, Theorem 2.1(1)] requires a curvature lower bound, the proof of it reveals that only the log-Sobolev inequality of the semigroup $\tilde{P}_{t}^{\mu}$ is used.
In order to prove Theorem 2.6, we need to derive a defective log-Sobolev inequality for the IPM of (2.13) uniformly in $\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d}):=\{\mu\in\mathscr{P}_{2}(\mathbb{R}^{d}),\mu(|\cdot|^{2})\leq M\}$ for any $M>0$. Together with tightness and uniform bounds on the densities of each IPM, this implies the result, as the next lemma shows.
@@
    Let $L^{\theta}=\frac{1}{2}\mathrm{tr}(a^{\theta}\nabla^{2})+\langle b^{\theta},\nabla\rangle$, $\theta\in\Theta$ be a family of conservative diffusion processes on $\mathbb{R}^{d}$ with associated semigroup $(P_{t}^{\theta})_{t\geq 0}$ satisfying the assumptions in (1) uniformly in $\theta$. Assume that the IPM $\bar{\mu}^{\theta}$ of $(P_{t}^{\theta})_{t\geq 0}$ form a tight subset of $\mathscr{P}(\mathbb{R}^{d})$ and that for any $R>0$, there exist positive constants $C_{1}(R)\leq C_{2}(R)$ such that $\bar{\rho}_{\theta}=\frac{\text{\rm{d}}\bar{\mu}^{\theta}}{\text{\rm{d}}x},\ \ \theta\in\Theta$ satisfy
@@
[…]
@@
(1) follows from he proof of [28, Theorem 2.1(1)] so that we omit it. It remains to prove (2). Observe that for any $R>0$, the classical Poincaré inequality for Lebesgue measure holds, i.e.
@@
$=\inf_{c\in\mathbb{R}}\frac{|B_{R}|}{\bar{\mu}^{\theta}(B_{R})}\int_{\mathbb{R}^{d}}(f-c)^{2}\frac{1_{B_{R}}}{|B_{R}|}\rho\text{\rm{d}}x$
@@
(2.19) $\bar{\mu}^{\theta}(f^{2})\leq\alpha(r_{0})\bar{\mu}^{\theta}(|\nabla f|^{2})+r_{0}\|f\|_{\infty}^{2},\ \ f\in C_{0}^{\infty}(\mathbb{R}^{d}),\bar{\mu}^{\theta}(f)=0,\theta\in\Theta.$
Observe that due to [38, Theorem 3.3.1], (1) implies the defective Poincaré inequality
$\bar{\mu}^{\theta}(f^{2})\leq C_{1}\bar{\mu}^{\theta}(|\nabla f|^{2})+C_{2}\bar{\mu}^{\theta}(|f|)^{2},\ \ \theta\in\Theta.$
By [35, Proposition 4.1.2(2)], we conclude from this and (2.19) that the Poincaré inequality holds:
@@
This together with the defective log-Sobolev inequality implies (2.17), see for instance the proof of [37, Corollary 1.3]. ∎
@@
Assume (H). Then the following assertions hold.
  1. $(1)$
    For any $\mu\in\mathscr{P}_{2}(\mathbb{R}^{d})$, (2.13) has a unique invariant probability measure denoted by $\Phi(\mu)$. Moreover, if $K_{I}<K_{2}$, there exists an $M>0$ such that $\Phi$ is a map on $\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d})$, where
    Moreover, there exists a constant $\varepsilon>0$ and $C_{1}>0$ such that
    (2.20) $\{\Phi(\mu)\}(\text{\rm{e}}^{\varepsilon|\cdot|^{2}})\leq C_{1},\ \ \mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d}).$
    Consequently, there exist constants $t_{0}>0,p>1,C_{2}>0$ such that
  2. $(2)$
    If in addition, $\sigma(x,\gamma)=\sigma(\gamma)$, then $\Phi(\mu)$ satisfies the log-Sobolev inequality
    (2.21) $\Phi(\mu)(f^{2}\log f^{2})\leq C_{L}\Phi(\mu)(|\nabla f|^{2}),$
    for some constant $C_{L}>0,$ any $\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d}),$ and $f\in C_{0}^{\infty}(\mathbb{R}^{d})$ with $\Phi(\mu)(f^{2})=1$.

### v4

@@ 2.2 Time-homogeneous decoupled SDEs
In this paper, we aim to prove exponential ergodicity of the time-homogeneous version (2.20) in $\mathbb{W}_{2}$ and relative entropy under the partially dissipative condition (H).
To this end, we first prove a log-Sobolev inequality for the invariant probability measure $\Phi(\mu)$ of the time-homogeneous decoupled SDEs, i.e. for any $\mu\in\mathscr{P}_{2}(\mathbb{R}^{d})$ we consider
We should remark that the log-Sobolev inequality for $\Phi(\mu)$ is uniform in $\mu$ with bounded second moments. We denote by $(\tilde{P}_{t}^{\mu})^{\ast}\nu$ the distribution of the solution to (2.23) with initial distribution $\nu\in\mathscr{P}_{2}(\mathbb{R}^{d})$ and associated time-homogeneous semigroup
$\tilde{P}_{t}^{\mu}f(x)=\int_{\mathbb{R}^{d}}f\text{\rm{d}}\big((\tilde{P}_{t}^{\mu})^{\ast}\delta_{x}\big),\ \ f\in\mathscr{B}_{b}(\mathbb{R}^{d}),x\in\mathbb{R}^{d},t\geq 0.$
2.2.1 Invariant probability measure
Theorem 2.6.
Assume (H). Then the following assertions hold.
  1. $(1)$
    For any $\mu\in\mathscr{P}_{2}(\mathbb{R}^{d})$, (2.23) has a unique invariant probability measure (IPM) denoted by $\Phi(\mu)$. Moreover, if $K_{I}<K_{2}$, there exists an $M>0$ such that $\Phi$ is a map on $\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d})$, where
  2. $(2)$
    There exist constants $\varepsilon>0$ and $C_{1}>0$ such that
    (2.24) $(\Phi(\mu))(\text{\rm{e}}^{\varepsilon|\cdot|^{2}})\leq C_{1},\ \ \mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d}).$
Remark 2.7.
We give some comments on the existence and uniqueness for invariant probability measure in Theorem 2.6(1). Under (H), when $\sigma=\sigma(\mu)$ and $b$ is locally Lipschitz continuous in spatial variable, one may use reflection couplings as in [17, Corollary 3, Lemma 1] to derive the exponential ergodicty in $L^{1}$-Wasserstein distance, see also [8, Theorem 13] for the multiplicative noise case where $\sigma$ and $b$ are assumed to be locally Lipschitz continuous in spatial variable. We should remark that in Theorem 2.6(1) above, the locally Lipschitz continuity of $b$ and $\sigma$ in spatial variable is not required. Another alternative method to derive the existence and uniqueness for invariant probability measure is to apply [16, Theorem 5.2(c)], which claims that for a continuous time Markov process with the extended generator $\tilde{L}$, if there exists $V\geq 1$ such that for some petite set $B$, some constants $c>0,b_{0}>0$,
$\tilde{L}V\leq-cV+b_{0}1_{B},$
then it is $V$-uniformly ergodic. Observe that the existence of petite set is the critical point in [16, Theorem 5.2(c)] and the level set $\{x:V(x)\leq n\}$ for some $n\geq 1$ may be the candidate.
Proof.
(1) Taking $y={\bf 0}$ in ((H)), we get
So, it follows from (2.22) and (2.2.1) that
for some constant $C_{0}>0$. Whence, an application of [55, Lemma 2.7(1)], along with (2.22) and (2.26), yields that (2.23) has an IPM.
By [51, Theorem 1.1(2)] and (H), one has Wang’s Harnack inequality, i.e. there exists $p_{0}>1$ such that for any $p>p_{0}$, one can find a constant $C>0$ such that
(2.27) $(\tilde{P}_{t}^{\mu}|f|)^{p}\leq\tilde{P}_{t}^{\mu}|f|^{p}\exp\left(C\frac{|x-y|^{2}}{1-\text{\rm{e}}^{-K_{1}t}}\right),\ \ f\in\mathscr{B}_{b}(\mathbb{R}^{d}),x,y\in\mathbb{R}^{d},t>0.$
So, [46, Theorem 1.4.1(3)] implies that (2.23) has a unique IPM written as $\Phi(\mu).$
Next, once $K_{I}<K_{2}$ is valid, [55, Lemma 2.8] implies that, for some constant $M>0,$ $\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d})$ is an invariant set of $\Phi$ so the second assertion in (1) holds true.
(2) Itô’s formula, (2.22) as well as (2.26) imply that for $\varepsilon=\frac{K_{2}}{4\delta_{1}}$ and $\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d}),$
$\text{\rm{d}}\text{\rm{e}}^{\varepsilon|\tilde{X}_{t}^{\mu}|^{2}}$ $\leq\varepsilon\text{\rm{e}}^{\varepsilon|\tilde{X}_{t}^{\mu}|^{2}}\left(-K_{2}|\tilde{X}_{t}^{\mu}|^{2}+2|\tilde{X}_{t}^{\mu}||b({\bf 0},\delta_{\bf 0})|+K_{I}\mu(|\cdot|^{2})+C_{0}+2\varepsilon\delta_{1}|\tilde{X}_{t}^{\mu}|^{2}\right)\text{\rm{d}}t$
$\leq$ $\varepsilon\text{\rm{e}}^{\varepsilon|\tilde{X}_{t}^{\mu}|^{2}}\left(-\frac{K_{2}}{4}|\tilde{X}_{t}^{\mu}|^{2}+\frac{4|b({\bf 0},\delta_{\bf 0})|^{2}}{K_{2}}+K_{I}M+C_{0}\right)\text{\rm{d}}t$
$\varepsilon\text{\rm{e}}^{\varepsilon|\tilde{X}_{t}^{\mu}|^{2}}\left(-\frac{K_{2}}{4}|\tilde{X}_{t}^{\mu}|^{2}+\frac{4|b({\bf 0},\delta_{\bf 0})|^{2}}{K_{2}}+K_{I}M+C_{0}\right)\leq c_{1}(\varepsilon)-c_{2}(\varepsilon)\text{\rm{e}}^{\varepsilon|\tilde{X}_{t}^{\mu}|^{2}}.$
Substituting this into (2.2.1) and taking $\tilde{X}_{0}^{\mu}={\bf 0}$ enables us to derive that for some constant $C_{1}>0$
This obviously implies (2.24) by a tightness argument.∎
Remark 2.8.
Our proof of the IPM’s uniqueness improves the approach in [55, Lemma 2.7(2)] by using Wang’s Harnack inequality [51, Theorem 1.1(2)]. This allows us to eliminate the need for the local Lipschitz of $b(\cdot,\mu),\sigma(\cdot,\mu)$ and the linear growth of $b(\cdot,\mu)$.
2.2.2 Wang’s Harnack inequality and hyperboundedness
To derive the hyperboundedness of $\tilde{P}_{t}^{\mu}$, a suitable Wang’s Harnack inequality is needed. In [53, Lemma 2.6], the author established a Harnack inequality for (2.23) in the case where the drift $b$ and diffusion coefficient $\sigma$ do not depend on the measure variable $\mu$. In the following theorem, we derive a uniform Wang’s Harnack inequality for (2.23) with respect to the measure parameter $\mu$. The proof is completely the same with that of [53, Lemma 2.6]. Since [53] is written in Chinese, for reader’s convenience, we will provide the proof. Note that (2.29) is different form (2.27), the factor $\text{\rm{e}}^{K_{2}t}-1$ in (2.29) is crucial for us to derive the hyperboundedness of $\tilde{P}_{t}^{\mu}$ below.
Lemma 2.9.
Assume (H). Then the following assertions hold.
[…]
(2.29) $(\tilde{P}_{t}^{\mu}|f|)^{p}(x)$ $\leq(\tilde{P}_{t}^{\mu}|f|^{p})(y)\exp\left((p-1)\left(h_{1}t+\frac{h_{2}K_{2}|x-y|^{2}}{\text{\rm{e}}^{K_{2}t}-1}\right)\right).$
(2) There exist constants $t_{0}>0,p>1,C_{2}>0$ such that
Proof.
Let $\xi_{t}=\frac{1}{K_{2}}(\text{\rm{e}}^{-K_{2}(t-t_{0})}-1),t\in[0,t_{0}],$ for fixed $t_{0}>0$. It is easy to see that
(2.30) $\xi_{t}^{\prime}+K_{2}\xi_{t}=-1,\ \ t\in[0,t_{0}].$
Consider the following coupling:
(2.31) $\begin{cases}\text{\rm{d}}\tilde{X}_{t}^{\mu}=b(\tilde{X}_{t}^{\mu},\mu)\text{\rm{d}}t+\sigma(\tilde{X}_{t}^{\mu},\mu)\text{\rm{d}}W_{t},\ \ \tilde{X}_{0}^{\mu}=x,\\ \text{\rm{d}}Y_{t}=b(Y_{t},\mu)\text{\rm{d}}t+\sigma(Y_{t},\mu)\Big(\text{\rm{d}}W_{t}+\hat{\sigma}(\tilde{X}_{t}^{\mu},\mu)\frac{\tilde{X}_{t}^{\mu}-Y_{t}}{\xi_{t}}\text{\rm{d}}t\Big),\ \ Y_{0}=y,t\in[0,t_{0}),\end{cases}$
where $\hat{\sigma}(x,\mu):=[\sigma^{\ast}(\sigma\sigma^{\ast})^{-1}](x,\mu)$. Set
$\text{\rm{d}}\tilde{W}_{t}:=\text{\rm{d}}W_{t}+\hat{\sigma}(\tilde{X}_{t}^{\mu},\mu)\frac{\tilde{X}_{t}^{\mu}-Y_{t}}{\xi_{t}}\text{\rm{d}}t.$
Obviously, (2.22) yields that
(2.32) $\left|\hat{\sigma}(\tilde{X}_{t}^{\mu},\mu)\frac{\tilde{X}_{t}^{\mu}-Y_{t}}{\xi_{t}}\right|^{2}\leq\delta_{2}^{-1}\frac{|\tilde{X}_{t}^{\mu}-Y_{t}|^{2}}{\xi_{t}^{2}}.$
According to (H), as explained in the proof of [51, Theorem 1.1], $(\tilde{W}_{t})_{t\in[0,t_{0}]}$ is an $n$-dimensional Brownian motion under $\text{\rm{d}}\mathbb{Q}=R_{t_{0}}\text{\rm{d}}\mathbb{P}$, where
$R_{s}=\exp\left(\int_{0}^{s}-\left\langle\hat{\sigma}(\tilde{X}_{t}^{\mu},\mu)\frac{\tilde{X}_{t}^{\mu}-Y_{t}}{\xi_{t}},\text{\rm{d}}W_{t}\right\rangle-\frac{1}{2}\int_{0}^{s}\left|\hat{\sigma}(\tilde{X}_{t}^{\mu},\mu)\frac{\tilde{X}_{t}^{\mu}-Y_{t}}{\xi_{t}}\right|^{2}\text{\rm{d}}t\right)$
with $\{R_{s}\}_{s\in[0,t_{0}]}$ being a uniformly integrable martingale and $\mathbb{Q}$-a.s. $\tilde{X}_{t_{0}}^{\mu}=Y_{t_{0}}$. Notice that (2.31) can be reformulated as follows:
$\begin{cases}\text{\rm{d}}\tilde{X}_{t}^{\mu}=b(\tilde{X}_{t}^{\mu},\mu)\text{\rm{d}}t+\sigma(\tilde{X}_{t}^{\mu},\mu)\text{\rm{d}}\tilde{W}_{t}-\frac{\tilde{X}_{t}^{\mu}-Y_{t}}{\xi_{t}}\text{\rm{d}}t,\ \ \tilde{X}_{0}^{\mu}=x,\\ \text{\rm{d}}Y_{t}=b(Y_{t},\mu)\text{\rm{d}}t+\sigma(Y_{t},\mu)\text{\rm{d}}\tilde{W}_{t},\ \ Y_{0}=y.\end{cases}$
Observe that (H) implies
$2\langle b(x,\mu)-b(y,\mu),x-y\rangle+\|\sigma(x,\mu)-\sigma(y,\mu)\|_{HS}^{2}$
$\leq(K_{1}+K_{2})|x-y|^{2}1_{\{|x-y|\leq r_{0}\}}-K_{2}|x-y|^{2}$
$\leq(K_{1}+K_{2})r_{0}|x-y|-K_{2}|x-y|^{2}.$
By Itô’s formula, we derive from this and (2.30) that
$\text{\rm{d}}\frac{|\tilde{X}_{t}^{\mu}-Y_{t}|^{2}}{\xi_{t}}$ $=-\frac{(\xi_{t}^{\prime}+2)|\tilde{X}_{t}^{\mu}-Y_{t}|^{2}}{\xi_{t}^{2}}\text{\rm{d}}t+\frac{2}{\xi_{t}}\langle(\sigma(\tilde{X}_{t}^{\mu},\mu)-\sigma(Y_{t},\mu))\text{\rm{d}}\tilde{W}_{t},\tilde{X}_{t}^{\mu}-Y_{t}\rangle$
$\quad+\frac{1}{\xi_{t}}\big(2\langle b(\tilde{X}_{t}^{\mu},\mu)-b(Y_{t},\mu),\tilde{X}_{t}^{\mu}-Y_{t}\rangle+\|\sigma(\tilde{X}_{t}^{\mu},\mu)-\sigma(Y_{t},\mu)\|_{HS}^{2}\big)\text{\rm{d}}t$
$\leq-\frac{|\tilde{X}_{t}^{\mu}-Y_{t}|^{2}}{\xi_{t}^{2}}\text{\rm{d}}t+\frac{(K_{1}+K_{2})r_{0}|\tilde{X}_{t}^{\mu}-Y_{t}|}{\xi_{t}}\text{\rm{d}}t$
(2.33) $\qquad\quad+\frac{2\langle(\sigma(\tilde{X}_{t}^{\mu},\mu)-\sigma(Y_{t},\mu))\text{\rm{d}}\tilde{W}_{t},\tilde{X}_{t}^{\mu}-Y_{t}\rangle}{\xi_{t}}$
$\leq-\frac{1}{2}\frac{|\tilde{X}_{t}^{\mu}-Y_{t}|^{2}}{\xi_{t}^{2}}\text{\rm{d}}t+\frac{1}{2}(K_{1}+K_{2})^{2}r_{0}^{2}\text{\rm{d}}t$
$\qquad\quad+\frac{2\langle(\sigma(\tilde{X}_{t}^{\mu},\mu)-\sigma(Y_{t},\mu))\text{\rm{d}}\tilde{W}_{t},\tilde{X}_{t}^{\mu}-Y_{t}\rangle}{\xi_{t}},\ \ t<t_{0}.$
For any $r>0$, it follows from (2.2.2) that
$\mathbb{E}_{\mathbb{Q}}\exp\left(r\int_{0}^{t_{0}}\frac{|\tilde{X}_{t}^{\mu}-Y_{t}|^{2}}{\xi_{t}^{2}}\text{\rm{d}}t\right)$
$\leq\exp\left(2r\frac{|x-y|^{2}}{\xi_{0}}+r(K_{1}+K_{2})^{2}r_{0}^{2}t_{0}\right)$
$\qquad\quad\times\mathbb{E}_{\mathbb{Q}}\exp\left(r\int_{0}^{t_{0}}\frac{4\langle(\sigma(\tilde{X}_{t}^{\mu},\mu)-\sigma(Y_{t},\mu))\text{\rm{d}}\tilde{W}_{t},\tilde{X}_{t}^{\mu}-Y_{t}\rangle}{\xi_{t}}\right)$
$\leq\exp\left(2r\frac{|x-y|^{2}}{\xi_{0}}+r(K_{1}+K_{2})^{2}r_{0}^{2}t_{0}\right)$
$\qquad\quad\times\left(\mathbb{E}_{\mathbb{Q}}\exp\left(128r^{2}\delta_{1}\int_{0}^{t_{0}}\frac{|\tilde{X}_{t}^{\mu}-Y_{t}|^{2}}{\xi_{t}^{2}}\right)\right)^{\frac{1}{2}}.$
where in the last step, we used $\|\sigma(x_{1},\mu)^{\ast}-\sigma(x_{2},\mu)^{\ast}\|^{2}\leq 4\delta_{1}$ and that for a continuous exponential integrable martingale $M_{t}$, it holds
$\mathbb{E}\text{\rm{e}}^{M_{t}}\leq(\mathbb{E}\text{\rm{e}}^{2\langle M\rangle_{t}})^{\frac{1}{2}}.$
Taking $r=\frac{1}{128\delta_{1}}$ and using $\xi_{0}=\frac{1}{K_{2}}(\text{\rm{e}}^{K_{2}t_{0}}-1)$, we arrive at
$\mathbb{E}_{\mathbb{Q}}\exp\left(\frac{1}{128\delta_{1}}\int_{0}^{t_{0}}\frac{|\tilde{X}_{t}^{\mu}-Y_{t}|^{2}}{\xi_{t}^{2}}\text{\rm{d}}t\right)$
(2.34) $\leq\exp\left(\frac{1}{32\delta_{1}}\frac{K_{2}|x-y|^{2}}{\text{\rm{e}}^{K_{2}t_{0}}-1}+\frac{1}{64\delta_{1}}(K_{1}+K_{2})^{2}r_{0}^{2}t_{0}\right).$
Let $\eta_{t}=-\hat{\sigma}(\tilde{X}_{t}^{\mu},\mu)\frac{\tilde{X}_{t}^{\mu}-Y_{t}}{\xi_{t}}$. Observe that for any $p>1$, it holds
$\mathbb{E}R_{t_{0}}^{\frac{p}{p-1}}=\mathbb{E}_{\mathbb{Q}}R_{t_{0}}^{\frac{1}{p-1}}$ $=\mathbb{E}_{\mathbb{Q}}\Bigg(\exp\left((p-1)^{-1}\int_{0}^{t_{0}}\langle\eta_{t},\text{\rm{d}}\tilde{W}_{t}\rangle-(p-1)^{-2}\int_{0}^{t_{0}}|\eta_{t}|^{2}\text{\rm{d}}t\right)$
$\qquad\quad\times\exp\left(\left(\frac{1}{2}(p-1)^{-1}+(p-1)^{-2}\right)\int_{0}^{t_{0}}|\eta_{t}|^{2}\text{\rm{d}}t\right)\Bigg)$
$\leq\left(\mathbb{E}_{\mathbb{Q}}\exp\left(\left((p-1)^{-1}+2(p-1)^{-2}\right)\int_{0}^{t_{0}}|\eta_{t}|^{2}\text{\rm{d}}t\right)\right)^{\frac{1}{2}}.$
Take $(p_{0}-1)^{-1}=\frac{1}{2}\wedge\frac{1}{256\delta_{2}^{-1}\delta_{1}}$. Then for any $p\geq p_{0}$, it holds
$(p-1)^{-1}+2(p-1)^{-2}\leq(p_{0}-1)^{-1}+2(p_{0}-1)^{-2}<2(p_{0}-1)^{-1}\leq\frac{1}{128\delta_{2}^{-1}\delta_{1}}.$
So, for any $p\geq p_{0}$, we derive from (2.32) and (2.2.2) that
(2.35) $\mathbb{E}R_{t_{0}}^{\frac{p}{p-1}}\leq\exp\left(\frac{1}{64\delta_{1}}\frac{K_{2}|x-y|^{2}}{\text{\rm{e}}^{K_{2}t_{0}}-1}+\frac{1}{128\delta_{1}}(K_{1}+K_{2})^{2}r_{0}^{2}t_{0}\right).$
Furthermore, it follows from the fact $\mathbb{Q}$-a.s. $\tilde{X}_{t_{0}}^{\mu}=Y_{t_{0}}$ and the Hölder inequality that for any $p>1$,
$(\tilde{P}_{t_{0}}^{\mu}|f|(y))^{p}=(\mathbb{E}_{\mathbb{Q}}(|f|(Y_{t_{0}})))^{p}=(\mathbb{E}_{\mathbb{Q}}(|f|(\tilde{X}_{t}^{\mu})))^{p}\leq\tilde{P}_{t_{0}}^{\mu}|f|^{p}(x)\left(\mathbb{E}R_{t_{0}}^{\frac{p}{p-1}}\right)^{p-1}.$
Combining this with (2.35), we finish the proof.
(2) Observing (1) and following the strategy adopted in [51, Corollary 1.3], it is more or less standard to derive from (2.29) that for some locally bounded function $(0,\infty)\ni t\mapsto C_{t}\in[0,\infty)$,
$\sup_{\{\Phi(\mu)\}(|f|^{p})\leq 1}(\tilde{P}_{t}^{\mu}|f|)^{2p}(x)$ $\leq C_{t}\exp\left(\frac{h_{2}K_{2}}{\text{\rm{e}}^{K_{2}t}-1}|x|^{2}\right),\ \ \mu\in\mathscr{P}_{2}(\mathbb{R}^{d}),x\in\mathbb{R}^{d}.$
(2.36) $\|\tilde{P}_{t_{0}}^{\mu}\|_{L^{p}(\Phi(\mu))\rightarrow L^{2p}(\Phi(\mu))}^{2p}\leq C_{t_{0}}\big(\Phi(\mu)\big)(\text{\rm{e}}^{\varepsilon|\cdot|^{2}})\leq C_{t_{0}}C_{1},\ \ \mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d}).$
The proof is completed. ∎
2.2.3 Log-Sobolev inequality
Before we prove the uniform log-Sobolev inequality for $\Phi(\mu)$ in $\mu$ with bounded second moments stated in Theorem 2.11, we give a general result Lemma 2.10. The first assertion in Lemma 2.10 is known by [39, Theorem 2.1(1)], which states that the hyperboundedness together with log-Sobolev inequality of the semigroup implies the defective log-Sobolev inequality for the corresponding IPM. Let us remark that although [39, Theorem 2.1(1)] requires a curvature lower bound, the proof of it reveals that only the log-Sobolev inequality of the semigroup is used. Lemma 2.10(2) is devoted to derive uniform log-Sobolev inequality from uniform defective log-Sobolev inequality for IPMs associated to a family of diffusion processes. The tightness of IPMs and uniform local bounds on the densities of each IPM will play an important role.
@@ 2. $(2)$
    Let $L^{\theta}=\frac{1}{2}\mathrm{tr}(a^{\theta}\nabla^{2})+\langle b^{\theta},\nabla\rangle$, $\theta\in\Theta$ be a family of conservative diffusion processes on $\mathbb{R}^{d}$ with associated semigroup $(P_{t}^{\theta})_{t\geq 0}$ and IPM $\bar{\mu}^{\theta}$ satisfying the assumptions in (1) uniformly in $\theta$. Assume that $\{\bar{\mu}^{\theta}\}_{\theta\in\Theta}$ form a tight subset of $\mathscr{P}(\mathbb{R}^{d})$ and that for any $R>0$, there exist positive constants $C_{1}(R)\leq C_{2}(R)$ such that $\bar{\rho}_{\theta}=\frac{\text{\rm{d}}\bar{\mu}^{\theta}}{\text{\rm{d}}x},\ \ \theta\in\Theta$ satisfy
@@
    (2.38) $\bar{\mu}^{\theta}(f\log f)-\bar{\mu}^{\theta}(f)\log\bar{\mu}^{\theta}(f)\leq C_{LS}\bar{\mu}^{\theta}(|\nabla f^{\frac{1}{2}}|^{2}),\ \ 0<f\in C_{0}^{\infty}(\mathbb{R}^{d}),\theta\in\Theta.$
@@
(1) follows from the proof of [39, Theorem 2.1(1)] so that we omit it. It remains to prove (2). Observe that for any $R>0$, the classical Poincaré inequality for Lebesgue measure holds, i.e.
@@
$=\inf_{c\in\mathbb{R}}\frac{|B_{R}|}{\bar{\mu}^{\theta}(B_{R})}\int_{\mathbb{R}^{d}}(f-c)^{2}\frac{1_{B_{R}}}{|B_{R}|}\bar{\rho}_{\theta}\text{\rm{d}}x$
@@
Observe that due to [49, Theorem 3.3.1], the uniform defective log-Sobolev inequality in $\theta$ due to (1) implies the defective Poincaré inequality
(2.40) $\bar{\mu}^{\theta}(f^{2})\leq C_{1}\bar{\mu}^{\theta}(|\nabla f|^{2})+C_{2}\bar{\mu}^{\theta}(|f|)^{2},\ \ f\in C_{0}^{\infty}(\mathbb{R}^{d}),\theta\in\Theta.$
@@
$\bar{\mu}^{\theta}(f^{2})\leq\alpha(r_{0})\bar{\mu}^{\theta}(|\nabla f|^{2})+r_{0}\|f\|_{\infty}^{2},\ \ f\in C_{0}^{\infty}(\mathbb{R}^{d}),\bar{\mu}^{\theta}(f)=0,\theta\in\Theta.$
By [49, Proposition 4.1.2(2)], we conclude from this and (2.40) that the Poincaré inequality holds:
@@
This together with the defective log-Sobolev inequality implies (2.38), see for instance the proof of [48, Corollary 1.3]. ∎
@@
Assume (H) and $\sigma(x,\gamma)=\sigma(\gamma)$. Then $\Phi(\mu)$ satisfies the log-Sobolev inequality
(2.41) $\Phi(\mu)(f\log f)-\Phi(\mu)(f)\log\Phi(\mu)(f)\leq C_{L}\Phi(\mu)(|\nabla f^{\frac{1}{2}}|^{2}),\ \ \mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d})$
for some constant $C_{L}>0$ and $f\in C_{0}^{\infty}(\mathbb{R}^{d})$ with $f>0$.
Remark 2.12.
In the symmetric case on $\mathbb{R}^{d}$, for $\mu(\text{\rm{d}}x)=\text{\rm{e}}^{-V}\text{\rm{d}}x/\int_{\mathbb{R}^{d}}\text{\rm{e}}^{-V}\text{\rm{d}}x$, the Dirichlet form is $\mathscr{E}(f,f)=\mu(|\nabla f|^{2})$. [7, Theorem 3.17] uses the Lyapunov conditions to derive the super Poincaré inequality and hence defective log-Sobolev inequality. The main idea of the proof is first proving a local super Poincaré inequality and then use Lyapunov condition to obtain a super Poincaré inequality on the whole space. This method relies crucially on the explicit information of $V$. In Theorem 2.11, we only present some conditions on the coefficients while the invariant probability measure is implicit.